# Effect of interleaflet friction on the dynamics of molecular rotors in lipid membranes

### v1

@@ Abstract
Molecular rotors form twisted conformations upon photoexcitation, with their fluorescent relaxation time serving as a measure of viscosity. They have been used to assess membrane viscosities but yield higher values compared to other methods. Here, we show that the rotor’s relaxation time is influenced by a combination of membrane viscosity and interleaflet friction. We present a theory for the relaxation time and obtain a correction factor that accounts for the discrepancy. If the membrane’s viscosity is known, molecular rotors may enable the extraction of the elusive interleaflet friction.
@@
Biological membranes encase cells and subcellular structures in living organisms, serving as barriers that regulate macromolecule transport, cell adhesion, mechanotransduction, and communication [1]. Biological membranes also facilitate enzymatic and metabolic activities vital for cellular processes. These critical functions of membranes are tightly dependent on their biophysical characteristics. However, despite extensive research, much remains to be discovered about important biophysical properties such as viscosity [2, 3, 4, 5, 6, 7, 8] and interleaflet friction [9].
Conventional rheometry techniques are not convenient to measure the viscosity of a membrane bilayer, especially at in vivo length scales. Instead, more intricate, microscopic methods are used. One such method is Fluorescence Recovery After Photobleaching (FRAP) [10] — lipids are marked by a fluorescent dye, a small area is photobleached, and the recovery of the fluorescence in the photobleached section is followed in time. Analyzing the fluorescence recovery kinetics, yields the Brownian diffusion coefficient ($D_{T}$) of the lipids, which is related to lipid translational resistance ($\Lambda$) through the Einstein relationship $D_{T}=k_{B}T/\Lambda_{T}$. For the simplest configurations, the membrane viscosity is inferred from resistance using the Saffman-Delbrück approximation [11], $\Lambda_{T}=4\pi\eta_{m}/\left[\ln\left(\frac{\eta_{m}}{\eta R}\right)-\gamma\right]$, where $\eta_{m}$ is the two-dimensional (2D) surface viscosity of the membrane, which can be related to a thin film viscosity $\eta_{m}^{*}$ as $\eta_{m}=\eta_{m}^{*}h$ [12]; $h$ is the thickness of the membrane, $\eta$ is the three-dimensional (3D) viscosity of the surrounding fluid, $R$ is the size of the diffusing particle, and $\gamma\simeq 0.5772$ is Euler’s constant. This formula for $D_{T}$ assumes $\frac{\eta_{m}}{\eta R}\gg 1$ so that $D_{T}>0$.
@@
Membrane viscosities obtained via fluorescence lifetimes of molecular rotors are, however, consistently larger than those independently obtained through diffusion measurements utilizing the Saffman-Delbrück formula [17, 22, 15, 14]. Although several potential reasons have been proposed to explain this discrepancy, no quantitative explanations are available. Here we provide a theoretical basis for this discrepancy by accounting for differences between rotor twist relaxation kinetics during calibration and during measurements within a membrane. Notably, we show that when embedded in membranes, molecular rotors measure a combination of the two-dimensional (2D) membrane viscosity and another difficult-to-measure quantity, interleaflet friction [23, 24, 25].
Current understanding of the position of the molecular rotor within the bilayer (e.g., [17]; see also Fig. 1) has it roughly spanning the mid-plane, which implies that when the molecule twists it induces shear between the two lipid layers in addition to rotational flows in each leaflet. With this picture in mind, we provide a theoretical prediction for the relaxation time of an initially twisted molecule as a function of both the membrane viscosity and interleaflet friction. We compare our theory to results given in the literature and show that the theoretical predictions can explain the discrepancy in viscosity measurements. As an additional outcome, it may be possible to extract interleaflet friction from measurements with molecular rotors if the membrane’s viscosity is known by other means, such as from FRAP or Fluorescence Correlation Spectroscopy (FCS) measurements. To proceed, we first present the problem of two counter-rotating disks in a membrane and find the typical relaxation time of an initially twisted molecular rotor. Then, we discuss the results and use them to reinterpret existing experiments in the literature.
@@
In order to determine the relaxation time of an excited molecular rotor in a membrane we will make a considerable simplification: we assume the molecular rotor is made of two counter-rotating disks, one in each leaflet (see Fig. 1). As outlined below the effective velocity of two counter-rotating disks can be interpreted as a combination of two axillary problems discussed below: 1) the velocity due to a disk rotating in a viscous 2D flow configuration following the original model of Saffman and Delbrück [11], and 2) the flow due to a disk rotating in a 2D “Brinkman fluid” [12], where there is additional friction on the leaflet, as in the case of a supported bilayer [26, 27, 28].
@@
Auxiliary Problems. Consider a disk of radius $R$ rotating with an angular velocity $\Omega$ in a 2D viscous fluid of viscosity $\eta_{m}$. The velocity field $\mathbf{v}$ in the membrane is governed by the Stokes equations, $\eta_{m}\nabla^{2}\mathbf{v}=\nabla p$, where $p$ is the pressure field. Here we have neglected the influence of the surrounding fluid on the membrane ($h\eta/\eta_{m}\ll 1$). From symmetry, we can assume that the flow field is in the $\mathbf{e}_{\theta}$ direction and is a function of $r$ alone, such that $\mathbf{v}=v(r)\mathbf{e}_{\theta}$. For a solution of this form the incompressibility requirement is implicitly satisfied and there are no gradients of pressure. The equation of motion is thus the Laplace equation:
@@
Next consider a disk that is rotating with angular velocity $\Omega$ in a 2D Brinkman fluid, such as for the case of a membrane close to a rigid wall. This flow model introduces an effective force on the membrane due to viscous stresses from the surrounding fluid. We follow Evans and Sackmann [12] in writing the equation of motion,
@@
Two counter-rotating disks in a membrane. Now consider a molecular rotor composed of two connected disks each of radius $R$ immersed between two membrane leaflets. The top disk rotates with angular velocity $\Omega_{1}$ and the bottom disk rotates in the opposite direction with angular velocity $\Omega_{2}$ (see Fig. 3). Suppose that the interleaflet friction is $\beta_{m}$ and as before assume that $\beta_{m}$ and $\eta_{m}$ are large enough to neglect the stress coming from the outer fluid. The equations of motion for such a case can be written as,
@@
The relative velocity satisfies a governing equation similar to the Brinkman, case Eq. (7), with the solution,
@@
where we used the fact that the relative angular velocity of the two disks is equal to the time-rate-of-change of the change of twist in the molecule, $2\Omega=\dot{\theta}$. The effective rotational resistance of the molecule can be obtained by computing the ratio of the net hydrodynamic torque to the relative angular velocity $\dot{\theta}$ as,
@@
Figure 4: Comparison with experiments. (a) Membrane viscosity evaluated from molecular diffusivity measurements ($\eta_{m}^{*}$) is consistently lower than those obtained from fluorescence lifetime measurements ($\eta^{*}_{m,\rm FLIM}$). Data from DOPC vesicles reported by Dent et al. [17] at different temperatures, and by Wu et al. [15] for different cholesterol concentrations at room temperature. Reported rotor diffusion measurements were obtained via Fluorescence Correlation Spectroscopy (FCS) or Molecular Dynamics (MD). (b) The fold differences between $\eta^{*}_{m,\rm FLIM}$ and $\eta_{m}^{*}$ are in alignment with those expected from Eq. (17). The values of $\beta_{m}$ used for the calculations are shown on the far right. $\beta_{m}$ at 293K was obtained from Ref. [29]. The same value was used at different cholesterol concentrations as $\beta_{m}$ is known to have a weak dependence on cholesterol concentration [23]. As reliable data for $\beta_{m}$ in DOPC vesicles was not available at other temperatures, $\kappa$ was assumed to remain the same across temperature. A * is used to indicate the $\beta_{m}$ satisfying this assumption. This choice is reasonable as $\kappa\propto\sqrt{\beta_{m}/\eta_{m}}$ should only have a weak temperature dependence as both $\beta_{m}$ and $\eta_{m}$ fall with temperature. See SI for additional details.
Fluorescence lifetime of a molecular rotor in a membrane. Following Förster and Hoffmann [18], we neglect inertia of the molecule and express the angular relaxation of the rotor through a spring-dashpot-like response, $\Lambda_{R}\dot{\theta}+k(\theta-\theta_{0})=0$. Here the spring constant $k$ is governed by molecular-scale interactions that drive the molecule back to its equilibrium angular orientation ($\theta_{0}$). The solution yields a classical decaying exponential with a relaxation time constant, $t_{2D}^{*}$, given by
@@
Membrane viscosity obtained via Fluorescence Lifetime Imaging Microscopy (FLIM) measurements is consistently larger than those obtained through diffusivity measurements. This fact can be seen in Fig. 4a, where available data in the literature from DOPC vesicles, including results obtained at different temperatures and with varying cholesterol fractions, are plotted to show the systematic overestimation of membrane viscosity by FLIM. The discrepancy can be quantitatively explained solely from differences between rotor hydrodynamics during calibration and during measurements within a membrane.
@@
In literature, membrane viscosity is usually reported as a thin film viscosity ($\eta_{m}^{*}$) with dimensions of bulk viscosity [17, 15, 21]. $\eta_{m}^{*}$ is related to the 2D membrane viscosity $\eta_{m}$ as $\eta_{m}=\eta_{m}^{*}h$ [12], where $h$ is the thickness of the membrane. Making this substitution in Eq.(17), we see that simply equating fluorescence lifetimes measured on a membrane to those obtained from calibration experiments in the bulk can overestimate membrane thin film viscosity ($\eta_{m}^{*}$) by a factor of $\kappa hK_{2}(\kappa R)/4K_{1}(\kappa R)$.
@@
These results underscore two key takeaways. Calibration curves obtained via 3D viscosity measurements should be corrected by multiplying the lifetime by a factor of $\left[\kappa hK_{2}(\kappa R)/4K_{1}(\kappa R)\right]^{2/3}$ for directly obtaining the accurate membrane viscosity. Second, if membrane viscosity $\eta_{m}^{*}$ is independently available, e.g., FCS, FRAP or MD simulations, Eq. (17) provides a convenient way to infer interleaflet friction — a hard-to-measure quantity, particularly on curved liposomes and in vivo. In this case, the interleaflet friction can be numerically extracted by solving $\eta^{*}_{m,\rm FLIM}/\eta_{m}^{*}=\kappa hK_{2}(\kappa R)/4K_{1}(\kappa R)$, where $\eta^{*}_{m,\rm FLIM}$ is the uncorrected membrane viscosity (obtained conventionally from bulk viscosity calibrated fluorescence lifetimes). These results also suggest molecular rotors with a larger radius are better suited for measuring interleaflet friction, thus providing guidance on the development of rotors optimized for sensing interleaflet friction (see SI). With further investigation on more molecular rotors and lipid systems, the provided framework can dramatically expand the use of molecular rotors as valuable molecular rheometry probes.

### v2

# Hydrodynamics of molecular rotors in lipid membranes

@@ Abstract
Molecular rotors form twisted conformations upon photoexcitation, with their fluorescence relaxation time serving as a measure of viscosity. They have been used to assess membrane viscosities but yield higher values compared to other methods. Here, we show that the rotor’s relaxation time is influenced by a combination of membrane viscosity and interleaflet friction. We present a theory for the relaxation time and obtain a correction factor that accounts for the discrepancy. If the membrane’s viscosity is known, molecular rotors may enable the extraction of the elusive interleaflet friction.
@@
I Introduction
Biological membranes encase cells and subcellular structures in living organisms, serving as barriers that regulate macromolecule transport, cell adhesion, mechanotransduction, and communication [1]. Biological membranes also facilitate enzymatic and metabolic activities vital for cellular processes. These critical functions of membranes are dependent on their biophysical characteristics. However, despite extensive research, much remains to be discovered about important biophysical properties such as viscosity [2, 3, 4, 5, 6, 7, 8] and interleaflet friction [9].
Conventional rheometry techniques are not convenient to measure the viscosity of a membrane bilayer, especially at in vivo length scales. Instead, more intricate, microscopic methods are used. One such method is Fluorescence Recovery After Photobleaching (FRAP) [10] — lipids are marked by a fluorescent dye, a small area is photobleached, and the recovery of the fluorescence in the photobleached section is followed in time. Analyzing the fluorescence recovery kinetics, yields the Brownian diffusion coefficient ($D_{T}$) of the lipids, which is related to lipid translational resistance ($\Lambda$) through the Einstein relationship $D_{T}=k_{B}T/\Lambda_{T}$. For the simplest configurations, the membrane viscosity is inferred from resistance using the Saffman-Delbrück approximation [11], $\Lambda_{T}=4\pi\eta_{m}/\left[\ln\left(\frac{\eta_{m}}{\eta R}\right)-\gamma\right]$, where $\eta_{m}$ is the two-dimensional (2D) surface viscosity of the membrane, which can be related to a thin film viscosity $\eta_{m}^{*}$ as $\eta_{m}=\eta_{m}^{*}h$ [12]; $h$ is the thickness of the membrane, $\eta$ is the three-dimensional (3D) viscosity of the surrounding fluid, $R$ is the radius of the diffusing particle, and $\gamma\simeq 0.5772$ is Euler’s constant. This formula for $D_{T}$ assumes $\frac{\eta_{m}}{\eta R}\gg 1$ so that $D_{T}>0$.
@@ I Introduction
Membrane viscosities obtained via fluorescence lifetimes of molecular rotors are, however, consistently larger than those independently obtained through diffusion measurements utilizing the Saffman-Delbrück formula [18, 23, 16, 14]. Although several potential reasons have been proposed to explain this discrepancy, no quantitative resolutions are available. While molecular rotors are not expected to yield viscosity measurements comparable to those obtained from conventional rheometers, in 3D, molecular rotor relaxation kinetics are predicted well by continuum hydrodynamics theories [19], and they yield accurate viscosity measurements. By extension, we hypothesize that molecular rotors will yield accurate measurements of membrane viscosities in 2D, provided the hydrodynamics of molecular rotors within membranes is accurately described. In alignment with this hypothesis, we present a theoretical hydrodynamic basis to address the discrepancy in membrane viscosities measured by molecular rotors by accounting for differences between rotor relaxation kinetics during calibration and during measurements within a membrane. Notably, we show that when embedded in membranes, molecular rotors measure a combination of the two-dimensional (2D) membrane viscosity and another difficult-to-measure quantity, interleaflet friction [24, 25, 26].
Current understanding of the position of the molecular rotor within the bilayer (e.g., [18]; see also Fig. 1) has it roughly spanning the mid-plane, with the bulk of the counter-rotating moieties of the rotor differentially localized to each of the bilayer leaflets (see SI for additional details). This implies that when the molecule twists it induces shear between the two lipid layers in addition to rotational flows in each leaflet. With this picture in mind, we provide a theoretical prediction for the relaxation time of an initially twisted molecule as a function of both the membrane viscosity and interleaflet friction. We compare our theory to results given in the literature and show that the theoretical predictions can explain the discrepancy in viscosity measurements. As an additional outcome, it may be possible to extract interleaflet friction from measurements with molecular rotors if the membrane’s viscosity is known by other means, such as from FRAP or Fluorescence Correlation Spectroscopy (FCS) measurements. To proceed, we first present the problem of two counter-rotating disks in a membrane and find the typical relaxation time of an initially twisted molecular rotor. Then, we discuss the results and use them to reinterpret existing experiments in the literature.
@@
II Theory
In order to determine the relaxation time of an excited molecular rotor in a membrane we will make a considerable simplification: we assume the molecular rotor is made of two counter-rotating disks, one in each leaflet (see Fig. 1). As outlined below the effective velocity of two counter-rotating disks can be interpreted as a combination of two axillary problems discussed below: 1) the velocity due to a disk rotating in a viscous 2D flow configuration following the original model of Saffman and Delbrück [11], and 2) the flow due to a disk rotating in a 2D “Brinkman fluid” [12], where there is additional friction on the leaflet, as in the case of a supported bilayer [27, 28, 29]. The application of a continuum hydrodynamic theory to predict rotor dynamics is consistent with well-established theories including the Saffman-Delbrück model [11], Förster-Hoffmann model [19], and the Stokes-Einstein-Debye model [30], all of which invoke continuum hydrodynamic arguments down towards molecular scale to obtain hydrodynamic resistances, as we aim to do below (see SI for additional details).
@@ II Theory
II.0.1 Rotating disk in a two-dimensional viscous fluid
. Consider a disk of radius $R$ rotating with an angular velocity $\Omega$ in a 2D viscous fluid of viscosity $\eta_{m}$. The velocity field $\mathbf{v}$ in the membrane is governed by the Stokes equations, $\eta_{m}\nabla^{2}\mathbf{v}=\nabla p$, where $p$ is the pressure field. Here we have neglected the influence of the surrounding fluid on the membrane as $h\eta/\eta_{m}\ll 1$.For characteristic values of variables reported in Fig. 4, $h\eta/\eta_{m}$ is less than $3\times 10^{-2}$. Furthermore, in 2D, the velocity field generated by rotation falls off faster (varies as $1/r$, see Eq.4) than that generated by translation (varies as $\log r$) [11]. As a consequence, viscous dissipation is localized to the membrane and the influence of bulk fluid viscosity on rotational drag can be safely neglected. From symmetry, we can assume that the flow field is in the $\mathbf{e}_{\theta}$ direction and is a function of $r$ alone, such that $\mathbf{v}=v(r)\mathbf{e}_{\theta}$. For a solution of this form the incompressibility requirement is implicitly satisfied and there are no gradients of pressure. The equation of motion is thus the Laplace equation:
@@ II.0.1 Rotating disk in a two-dimensional viscous fluid
II.0.2 Rotating disk in a two-dimensional Brinkman fluid
. Next consider a disk that is rotating with angular velocity $\Omega$ in a 2D Brinkman fluid, such as for the case of a membrane close to a rigid wall. This flow model introduces an effective force on the membrane due to viscous stresses from the surrounding fluid. We follow Evans and Sackmann [12] in writing the equation of motion,
@@ II.0.2 Rotating disk in a two-dimensional Brinkman fluid
II.0.3 Two counter-rotating disks in a membrane
Now consider a molecular rotor composed of two connected disks each of radius $R$ immersed between two membrane leaflets. The top disk rotates with angular velocity $\Omega_{1}$ and the bottom disk rotates in the opposite direction with angular velocity $\Omega_{2}$ (see Fig. 3). Suppose that the interleaflet friction is $\beta_{m}$ and as before assume that $\beta_{m}$ and $\eta_{m}$ are large enough to neglect the stress coming from the outer fluid. The equations of motion for such a case can be written as,
@@ II.0.3 Two counter-rotating disks in a membrane
The relative velocity satisfies a governing equation similar to the Brinkman case, Eq. (7), with the solution,
@@
where we used the fact that the relative angular velocity of the two disks is equal to the time-rate-of-change of the change of twist in the molecule, $2\Omega=\dot{\theta}$. The above equation also reveals that molecular rotors, while undergoing small angular displacements of approximately $1\;nm$, can induce long range velocity disturbances that span several molecular radii across the membrane (see SI). The effective rotational resistance of the molecule can be obtained by computing the ratio of the net hydrodynamic torque to the relative angular velocity $\dot{\theta}$ as,
@@
Figure 4: Comparison with experiments. (a) Membrane viscosity evaluated from molecular translational diffusivity measurements ($\eta_{m}^{*}$) is consistently lower than those obtained from fluorescence lifetime measurements ($\eta^{*}_{m,\rm FLIM}$). Data from DOPC vesicles reported by Dent et al. [18] at different temperatures, and by Wu et al. [16] for different cholesterol concentrations at room temperature. Reported rotor diffusion measurements were obtained via Fluorescence Correlation Spectroscopy (FCS) or Molecular Dynamics (MD). (b) The fold differences between $\eta^{*}_{m,\rm FLIM}$ and $\eta_{m}^{*}$ are in alignment with those expected from Eq. (17). The values of $\beta_{m}$ used for the calculations are shown on the far right. $\beta_{m}$ at 293K was obtained from Ref. [31]. The same value was used at different cholesterol concentrations as $\beta_{m}$ is known to have a weak dependence on cholesterol concentration [24]. As reliable data for $\beta_{m}$ in DOPC vesicles was not available at other temperatures, $\kappa$ was assumed to remain the same across temperature. A * is used to indicate the $\beta_{m}$ satisfying this assumption. This choice is reasonable as $\kappa\propto\sqrt{\beta_{m}/\eta_{m}}$ should only have a weak temperature dependence as both $\beta_{m}$ and $\eta_{m}$ fall with temperature. See SI for additional details.
II.0.4 Fluorescence lifetime of a molecular rotor in a membrane
. Following Förster and Hoffmann [19], we neglect inertia of the molecule and express the angular relaxation of the rotor through a spring-dashpot-like response, $\Lambda_{R}\dot{\theta}+k(\theta-\theta_{0})=0$. Here the spring constant $k$ is governed by molecular-scale interactions that drive the molecule back to its equilibrium angular orientation ($\theta_{0}$). The solution yields a classical decaying exponential with a relaxation time constant, $t_{2D}^{*}$, given by
@@ II.0.4 Fluorescence lifetime of a molecular rotor in a membrane
III Discussion
Membrane viscosities obtained via Fluorescence Lifetime Imaging Microscopy (FLIM) measurements are consistently larger than those obtained through diffusivity measurements. This fact can be seen in Fig. 4a, where available data in the literature from DOPC vesicles, including results obtained at different temperatures and with varying cholesterol fractions, are plotted to show the systematic overestimation of membrane viscosity by FLIM. The discrepancy can be quantitatively explained solely from differences between rotor hydrodynamics during calibration and during measurements within a membrane.
@@ III Discussion
In the literature, membrane viscosity is usually reported as a thin film viscosity ($\eta_{m}^{*}$) with dimensions of bulk viscosity [18, 16, 22]. $\eta_{m}^{*}$ is related to the 2D membrane viscosity $\eta_{m}$ as $\eta_{m}=\eta_{m}^{*}h$ [12], where $h$ is the thickness of the membrane. Making this substitution in Eq. (17), we see that simply equating fluorescence lifetimes measured on a membrane to those obtained from calibration experiments in the bulk can overestimate membrane thin film viscosity ($\eta_{m}^{*}$) by a factor of $\kappa hK_{2}(\kappa R)/4K_{1}(\kappa R)$.
@@
These results underscore two key takeaways. Calibration curves obtained via 3D viscosity measurements should be corrected by multiplying the lifetime by a factor of $\left[\kappa hK_{2}(\kappa R)/4K_{1}(\kappa R)\right]^{2/3}$ for directly obtaining the accurate membrane viscosity. Second, if membrane viscosity $\eta_{m}^{*}$ is independently available, e.g., FCS, FRAP or MD simulations, Eq. (17) provides a convenient way to infer interleaflet friction — a hard-to-measure quantity, particularly on curved liposomes and in vivo. In this case, the interleaflet friction can be numerically extracted by solving $\eta^{*}_{m,\rm FLIM}/\eta_{m}^{*}=\kappa hK_{2}(\kappa R)/4K_{1}(\kappa R)$, where $\eta^{*}_{m,\rm FLIM}$ is the uncorrected membrane viscosity (obtained conventionally from bulk viscosity calibrated fluorescence lifetimes). These results also suggest that molecular rotors with a larger radius are better suited for measuring interleaflet friction, thus providing guidance on the development of rotors optimized for sensing interleaflet friction (see SI). With further investigation on more molecular rotors and lipid systems, the provided framework can expand the use of molecular rotors as valuable molecular rheometry probes.
Acknowledgements
We note that this work began in 2018 when NO was a postdoc with HAS, but stalled because the researchers were unable to see the connection of the theory to the experimental literature. A conversation between VCS and HAS in Spring 2024 led VCS to reconsider the findings and demonstrate the ability of the theory to rationalize the existing experimental measurements.